OpenStudy (anonymous):

HELP PLEASE ? algebra 1 Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x3y4)2

OpenStudy (anonymous):

hate algebra 1.. in it and it SUCKKKKS

OpenStudy (anonymous):

I'm a senior taking it for grade forgiveness :(

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

is this your expression?: \[\large (3x^3y^4)^2 \] ??

OpenStudy (anonymous):

Yes @dpaInc

OpenStudy (anonymous):

\[\large (3x^3y^4)^2=(3x^3y^4)(3x^3y^4)=3*x^3*y^4*3*x^3*y^4 \] rearrange them and simplify...

OpenStudy (anonymous):

That's what i don't get, how to explain it

OpenStudy (anonymous):

multiply: 3*3 = x^3 * x^3 = y^4 * y^4 =